# IEEEexample.bib

### v2

@@ III-D2 Segmentation Task
We used average precision (AP) to measure the accuracy of the methane plume segmentation task. The formula for calculating AP is as follows[he2018hashing]:
@@ 5. Calculate the AP value:
where $(R_{i+1}-R_{i})$ is the length of recall interval, and $P_{inter}(R_{i}+1)$ is the interpolated precision at recall $R_{i}+1$. The AP value is calculated by summing up all interpolated precision values weighted by the lengths of recall intervals between adjacent recalls.
@@ IV-A Simulation
To verify the accuracy of the simulated plumes, we consulted relevant studies[gorrono2023understanding]. The integrated methane emission (IME) values for a typical methane plume should exhibit a linear relationship with its emission rate. We computed the IME values for all 2D plume snapshots, plotted them against their respective emission rates in a scatter plot, and fitted a linear curve, from which we obtained the coefficient of determination. The findings show that the coefficient of determination for the linear curve was 0.88 (Fig. 5), indicating a linear correlation between the IME values and emission rates of these 2D plume snapshots. Therefore, we can conclude that the quality of simulated methane plumes is reasonable to a large extent.
@@
We calculated all 2D plume snapshots simulated by palm software and their IME values. We plotted a scatter diagram of IME values of plume snapshots and their corresponding emission rates and fitted a linear curve. The results show that there is a high correlation between IME values and emission rates of these plume snapshots.
We employ simulation techniques to create EnMAP datasets with plume signals. Despite the fact that the production process adheres to theoretical foundations, the physical model remains a simplified version of actual conditions. Thus, it is essential to confirm the validity of incorporating plume signals into the EnMAP base map. In the absence of actual labels, we utilize an end-to-end method to confirm the simulated EnMAP images. Firstly, a random 2D plume snapshot is captured, and a transmittance matrix is constructed following the procedure detailed in Section II. Next, a random EnMAP base map is chosen and combined with the transmittance matrix using the Hadamard product to generate a simulated EnMAP image. Following this, the Mag1c tool is employed to extract the methane concentration distribution map from the simulated EnMAP image. Finally, we compare the 2D plume snapshot to the methane concentration distribution map. If their methane concentration distributions match closely, the plume signal incorporation process can be deemed valid. In Fig. 6, the left image shows a 2D plume snapshot, whereas the right image presents the methane concentration distribution map derived by Mag1c from the relevant EnMAP data. It is notable that the right image contains some additional noise compared to the left, but the primary features are quite similar, thereby validating the plume signal incorporation process.
@@
Due to the absence of the ground truth value of methane concentration, we verify it through an end-to-end method: comparing the methane plume extracted from the simulated EnMAP image using mag1c with the methane plume added to the EnMAP base map. The results show that the methane plume extracted by mag1c closely resembles the plume incorporated into the EnMAP base map, indicating that the addition of the plume signal is appropriate.
@@ IV-B Methane Concentration Inversion
Since PRISMA and EnMAP have 49 and 41 bands remaining after removing the sensitive bands for water vapor and carbon dioxide, we first resampled the PRISMA data into 41 bands in the spectral direction, referring to the center wavelengths of the 41 bands of EnMAP that participate in prediction. Then we used the trained U-net to invert the methane plume in this region and compared it with the Mag1c inversion results based on 49 PRISMA bands. As shown in Fig. 7, there are three representative inversion visualization results. It can be seen that in some data, the inversion results of U-net and Mag1c are consistent, while there are some discrepancies in the inversion results of the two methods in some data. Due to the absence of a true concentration value for the methane plume column in the WENT region, it is impossible to ascertain which method yields the most accurate inversion results. We explore possible reasons for the disparities in the inversion outcomes between the two approaches: 1. The WENT region is uneven and near the ocean, and the complexity of the noise distribution in the collected PRISMA data affects the two methods differently. 2. The resampling process of the PRISMA data results in some loss of plume information, causing disparities in the inversion results of U-net and Mag1c. 3. There is a limited number of EnMAP base maps used for U-net training available in rugged areas, leading U-net to likely predict methane column concentrations based on smoother terrains.
@@
Sample illustrations of U-net’s generalization performance on PRISMA..
@@ IV-C Methane Plume Segmentation
We trained the Mask R-CNN model using two versions of the plume segmentation dataset described in Section II.C. One version used simulated hyperspectral images as input (hyperspectral image), while the other version used methane plume distribution maps (single). However, the labels for both versions of the dataset were identical. Similarly, we separately validated the models using validation sets corresponding to two versions of the dataset with identical labels. Note that the Active Contour method was only validated using single methane plume maps as input, since it is an unsupervised image segmentation method that needs to explicitly relate pixel values to task-dependent physical property. The results showed that the Mask R-CNN model achieved a higher precision than Active Contour (Table II). Nonetheless, when evaluating the AP metric at $AP_{95}$, the Active Contour method surpassed Mask R-CNN in accuracy, suggesting that Active Contour excels at isolating individual plume masks. Despite this, Active Contour exhibits lower overall performance compared to Mask R-CNN due to its inability to effectively distinguish overlapping plumes and its propensity to produce small extraneous segmented areas.
@@
Both versions of Mask R-CNN demonstrate superior precision in extracting the plume mask from the methane plume distribution map compared to extracting it directly from hyperspectral images. This is primarily due to two factors: first, the EnMAP data, produced through a two-step simulation process, inherently contain more errors than the one-step simulation method. Second, extraction of the EnMAP plume mask involves both inversion and segmentation steps, which increases the likelihood of inaccuracies compared to direct extraction from the methane plume distribution map. Regarding the ”single” Mask R-CNN, despite its capability to efficiently segment overlapping plumes, its segmentation outcome still exhibits some errors (Fig. 8). These errors can be generally categorized as follows: 1. If a plume is divided into two sections, the single Mask R-CNN may incorrectly segment it into two separate plumes. 2. For overlapping plumes, the single Mask R-CNN may generate redundant or duplicate segmentations.
@@
The region outlined by the blue border in the figure shows the segmentations of the methane plume produced by different methods.
To evaluate the Mask R-CNN model’s ability to generalize to real-world contexts, we utilized mag1c to reverse-engineer methane concentration maps from various PRISMA images within the WENT area. Subsequently, both Active Contour and Mask R-CNN were applied to delineate methane plumes. The blue boxes in Fig. 9 denote these identified plumes. It is evident that Active Contour maps all connected pixel regions, whereas Mask R-CNN typically highlights funnel-shaped pixel regions, indicating some resistance to noise. Despite this, both methods identified a comparable number of methane plumes. We acknowledge the absence of precise data on the exact count of methane emission sources (or plumes) in the WNET area. Nevertheless, visual inspection reveals that both techniques detected the major plumes and overlooked several smaller ones on the right. This omission could be attributed to the rugged coastal terrain of the WENT region, where plume dispersion is influenced by intricate factors like topography and sea-land breeze dynamics, resulting in more complex plume formations.
@@ IV-D Estimation of Methane Emission Rate
We trained ResNet-50 and AlexNet for emission rate estimation using the training dataset and verified the accuracy of the model estimation in the corresponding test set (Table III). The results show that ResNet-50 outperforms the IME algorithm in the validation set. The emission rate estimation accuracy of AlexNet is slightly higher than that of the IME algorithm, but lower than ResNet-50. This aligns with with the relevant experience in the field of deep learning, where a deeper network means that each layer of the network can perform less complex tasks, thus obtaining more nonlinear expressive capabilities, learning more complex mappings[2016On], and enhancing the generalization performance and robustness of the network.
@@
To assess the ResNet-50’s ability to generalize to real-world data, we employed Mag1c to invert methane concentration distribution maps derived from RPISMA data in the WENT region and manually identified 28 methane plumes. We estimated their emission rates using both the IME method and the pre-trained ResNet-50 model, constructed scatter plots, and performed linear curve fitting. As shown in Fig. 10, the scatter plot illustrates that the majority of points are aligned closely to a straight line, indicating that the predictions from both approaches are typically in good agreement with the real-world dataset. However, it is also apparent that the ResNet-50 model shows a tendency to saturate in areas with lower emission rates.
@@
We manually extracted 28 methane plumes from the inversion results of PRISMA’s WENT region and predicted their emission rates using the IME method and the trained ResNet-50 model separately. The prediction results are shown in the figure above. Both methods yielded comparable results. However, it can also be observed that ResNet-50 shows some saturation in regions with low emission rates.
@@ IV-E Serial Network
While individual models demonstrated high accuracy for each specific task, the efficacy of serializing multi-network models still warrants further examination. Initially, we combined Mask R-CNN and ResNet-50 to handle the tasks of segmenting methane plumes from concentration maps and estimating their emission rates. We assessed the accuracy of these emission rate estimates on the validation set under three scenarios: 1. For true positive segmented plumes, we used pre-trained ResNet-50 to estimate emission rates and compared them with the label’s emission rates. 2. For false positive segmented plumes, we similarly used ResNet-50 to estimate their emission rates and compared them against 0 kg/h. 3. For misdetected plumes, we assumed their emission rate estimates to be 0 kg/h and compared these estimates with the corresponding labels. Ultimately, RMSE and MAE between all predicted emission rates and label values were calculated (Table IV). The results indicated that the accuracy of emission rate estimation using Mask R-CNN + ResNet-50 was significantly lower than using ResNet-50 for single-task emission rate estimation. It indicates that the serialization of Mask R-CNN and ResNet-50 could introduce additional errors in overall performance of methane flux rate estimation.
@@
To further explore the additional error sources, we selected the segmented plumes with large errors (e.g. greater than 500 kg/h). The results revealed two sources of additional errors: 1. Errors caused by small patches that actually contain only noises; 2. Errors caused by false positive patches.
We also studied the distribution of small noise patches and found that most are characterized by a mask size of smaller than 300 pixels, whereas the real plume mask size is rarely smaller than 300 pixels. Therefore, we removed the segmented masks whose sizes are smaller than 300 pixels and revalidate the accuracy of Mask R-CNN + ResNet-50 for emission rate estimation.
@@ IV-F Multi-task Learning I (MTL-01)
We trained the $MTL-01$ network and after 30 epochs of training, $MTL-01$ achieved higher accuracy in the plume segmentation problem compared to training Mask R-CNN alone. Consequently, the loss in the estimation of the emission rate of multiple methane flows was also smaller than in the case of serialization of $Mask\,R-CNN$ and $ResNet-50$. Furthermore, we simulated a smaller dataset to train the networks using the two approaches and $MTL-01$ exhibited even greater accuracy compared to serializing $Mask\,R-CNN$ and $MTL-01$.
We can observe that the segmentation of additional plumes is suppressed(Fig. 9), thereby reducing the estimated emission rate. $MTL-01$ tends to segment the main parts of the jet stream, ignoring some smaller tail parts. This has brought improvements to both the plume segmentation task and the estimation of the emission rate(Table II, Table IV).
Figure 11: Comparison of serial networks and multi-task learning I (MTL-01).
@@ IV-G Multi-task Learning II (MTL-02)
We trained the MTL-02 model, which converged after 30 epochs, with both the loss function values and the validation accuracy values rapidly decreasing. Ultimately, MTL-02 achieved slightly higher accuracy in methane concentration inversion compared to U-net(Table I), and higher accuracy in plume segmentation compared to Hyper Mask R-CNN(Table II). This confirms the viewpoint that multi-task learning could improve the convergence speed and fitting accuracy of the model. This may be attributed to the shared information among multiple types of labels.
@@ IV-H Limitations
Monitoring methane plumes is an intricate issue, and the current research in this domain is neither thoroughly explored nor systematic. Our proposed method also has some constraints: 1. The shapes of the simulated plumes we used mostly adhere to a standard Gaussian model, which represents an ideal scenario. However, in reality, methane emissions from point sources are influenced by a variety of factors like topography and humidity, leading to non-Gaussian plumes. Whether our method applies to non-Gaussian plumes remains an open question; 2. The maximum overlap ratio of the plumes we examined is $15\%$. However, naturally, many methane emission sources are situated close to each other, and the overlap ratio of two plumes can be significantly higher than $15\%$. In such scenarios, even though we can generate highly overlapped methane plumes, it is challenging to provide precise training labels, necessitating the exploration of point-based semi-supervised learning techniques to achieve reasonable segmentation outcomes for these highly overlapped plumes; 3. The acquisition of hyperspectral satellite imagery is significantly affected by weather conditions, and hyperspectral image data sources are scarce in low-latitude coastal regions. Hence, a solution relying solely on hyperspectral images may not fulfill high-frequency monitoring needs. It is essential to devise an ensemble estimation method based on multiple data sources.
@@ V Conclusion
Currently, there are high uncertainties in fugitive methane emission monitoring methods using satellite imaging spectrometers. We have developed a demonstrative approach that utilizes deep learning models and simulated hyperspectral images for methane concentration inversion, plume segmentation, and emission rate estimation in landfills. Specifically, we employed U-net for methane concentration inversion, achieving higher accuracy and reduced noise compared to mag1c. We utilized Mask R-CNN for plume segmentation, avoiding subjectivity in the segmentation process. For methane plume emission rate estimation, ResNet-50 is employed to overcome the reliance on wind speed observations. We also discussed the additional errors of emission rate estimation introduced by serial networks and addressed this issue by training a multi-task model (MTL-01) to suppress the errors and improve the precision of plume segmentation. Additionally, a multi-task model (MTL-02) that simultaneously performs methane concentration inversion and plume segmentation is proposed, achieving higher validation accuracy than individual single-task models. Finally, we also indicated the limitations of the current approach, highlighting the need for future research into highly overlapped plumps, non-Gaussian plumes, and the synthesis of multi-source data for inversion.

### v3

@@ III-D2 Segmentation Task
Average precision (AP) is used to measure the accuracy of the methane plume segmentation task. The formula for calculating AP is as follows[he2018hashing]:
@@ 5. Calculate the AP value:
where $(R_{i+1}-R_{i})$ represents the span of the recall interval, and $P_{inter}(R_{i}+1)$ denotes the interpolated precision at a recall of $R_{i}+1$. The AP value is computed by aggregating all interpolated precision values, each weighted by their respective recall interval lengths between consecutive recalls.
@@ IV-A Simulation
To verify the accuracy of simulated plumes, we consulted relevant studies[gorrono2023understanding]. The integrated methane emission (IME) values for a typical methane plume should exhibit a linear relationship with its emission rate. We computed the IME values for all 2D plume snapshots, plotted them against respective emission rates in a scatter plot, and fitted a linear curve, from which the coefficient of determination is derived. The findings show that the coefficient of determination for the linear curve was 0.88 (Fig. 5), indicating a linear correlation between the IME values and emission rates of these 2D plume snapshots. Therefore, we conclude that the quality of simulated methane plumes is reasonable to a large extent.
@@
We analyzed all 2D plume images generated by the PALM software and determined their IME values. A scatter plot was created to display the IME values of the plume images against their corresponding emission rates, and a linear fit was applied. The findings indicate a strong correlation between the IME values and emission rates of these plume images.
We utilize simulation methods to create EnMAP datasets that include plume signals. Although the production process follows theoretical principles, the physical model is a simplified approximation of real-world conditions. Therefore, it is crucial to verify the accuracy of incorporating plume signals into the EnMAP base map. Without actual labels, we apply an end-to-end approach to validate the simulated EnMAP images. First, a random 2D plume snapshot is taken and a transmittance matrix is created following the method described in Section II. Next, a random EnMAP base map is chosen and combined with the transmittance matrix using the Hadamard product to generate a simulated EnMAP image. Following this, the Mag1c tool is employed to extract the methane concentration distribution map from the simulated EnMAP image. Finally, we compare the 2D plume snapshot with the methane concentration distribution map. If their methane concentration distributions match closely, the plume signal incorporation process can be deemed valid. In Fig. 6, the left image shows a 2D plume snapshot, while the right image presents the methane concentration distribution map derived by Mag1c from the relevant EnMAP data. It is notable that the right image contains some additional noise compared to the left image, but the primary features are quite similar, thereby validating the plume signal incorporation process.
@@
Due to the absence of the ground truth value of methane concentration, the evaluation is performed through an end-to-end method: comparing the methane plume extracted from the simulated EnMAP image using mag1c with the methane plume added to the EnMAP base map. The results show that the methane plume extracted by mag1c closely resembles the plume incorporated into the EnMAP base map, indicating that the addition of the plume signal is appropriate.
@@ IV-B Methane Concentration Inversion
After eliminating the sensitive bands for water vapor and carbon dioxide, PRISMA and EnMAP retained 49 and 41 bands, respectively. The PRISMA data were then resampled to 41 bands in the spectral dimension, aligned with the center wavelengths of the 41 EnMAP bands used in the prediction. Subsequently, the pre-trained U-net was applied to invert the methane plume in the area, and its results were compared to the Mag1c inversion based on the 49 PRISMA bands. As depicted in Fig. 7, three representative inversion visualizations were presented. It is observable that U-net and Mag1c provide consistent inversion results in certain datasets, while some differences are evident in others. Due to the lack of true methane plume concentration values in the WENT region, determining the most accurate method is challenging. We investigated potential reasons for inconsistencies between the two approaches: 1. The uneven terrain of the WENT region and proximity to the ocean create a complex noise distribution in the PRISMA data, affecting the methods differently. 2. The PRISMA data resampling process results in some plume information loss, leading to disparities in U-net and Mag1c inversion results. 3. The limited number of EnMAP base maps used for U-net training in rugged areas may cause U-net to predict methane column concentrations based on smoother terrain.
@@
Sample illustrations of U-net’s generalization performance on PRISMA image.
@@ IV-C Methane Plume Segmentation
The Mask R-CNN model was trained on two versions of the plume segmentation dataset as detailed in Section II.C. One version used simulated hyperspectral images (hyperspectral image) as input, while the other used methane plume distribution maps (single). Despite this, the labels for both versions of the data set were identical. Similarly, model validation was performed separately using validation sets corresponding to the two versions of the data set, maintaining identical labels. Note that the Active Contour method was validated exclusively using single methane plume maps as input because it is an unsupervised image segmentation technique that requires an explicit relationship between pixel values and task-dependent physical properties. The results indicated that the Mask R-CNN model achieved higher precision than the Active Contour method (Table II). However, when assessing the AP metric at $AP_{95}$, the Active Contour method outperformed Mask R-CNN, suggesting that Active Contour excels in isolating individual plume masks. Despite this, the overall performance of Active Contour is inferior to that of Mask R-CNN due to its difficulty in effectively distinguishing overlapping plumes and its tendency to create small, extraneous segmented areas.
@@
Both versions of Mask R-CNN demonstrate superior precision in extracting the plume mask from the methane plume distribution map compared to extracting it directly from hyperspectral images. This is primarily due to two factors: first, the EnMAP data, produced through a two-step simulation process, inherently contain more errors than the one-step simulation method. Second, extraction of the EnMAP plume mask involves both inversion and segmentation steps, which increases the likelihood of inaccuracies compared to direct extraction from the methane plume distribution map. Regarding the ”single” Mask R-CNN, despite its capability to efficiently segment overlapping plumes, its segmentation outcome still exhibits some errors (Fig. 8). These errors can generally be classified as follows: 1. If a plume is divided into two sections, the single Mask R-CNN may incorrectly segment it into two separate plumes. 2. For overlapping plumes, the single Mask R-CNN may generate redundant or duplicate segmentations.
@@
The region outlined by the blue border shows the segmentation of methane plumes produced by different methods.
To evaluate the Mask R-CNN model’s ability to generalize to real-world contexts, mag1c is used to reverse-engineer methane concentration maps from various PRISMA images within the WENT area. Subsequently, both Active Contour and Mask R-CNN were applied to delineate methane plumes. The blue boxes in Fig. 9 denote these identified plumes. It is evident that Active Contour maps all connected pixel regions, whereas Mask R-CNN typically highlights funnel-shaped pixel regions, indicating some resistance to noise. Despite this, both methods identified a comparable number of methane plumes. We acknowledge the absence of precise data on the exact count of methane emission sources (or plumes) in the WNET area. Nevertheless, visual inspection reveals that both techniques detected the major plumes and overlooked several smaller ones on the right. This omission could be attributed to the rugged coastal terrain of the WENT region, where plume dispersion is influenced by intricate factors like topography and sea-land breeze dynamics, resulting in more complex plume formations.
@@ IV-D Estimation of Methane Emission Rate
We employed ResNet-50 and AlexNet to estimate emission rates by training them on the provided dataset and evaluating the accuracy of their estimations on the corresponding test set (Table III). The results reveal that ResNet-50 outperforms the IME algorithm in the validation set. Although AlexNet’s estimation accuracy is slightly higher than that of the IME algorithm, it still does not match ResNet-50’s performance. This aligns with the common understanding in deep learning that deeper networks enable each layer to address simpler tasks. As a result, more complex nonlinear representations are possible, which helps to learn intricate mappings[2016On], and boosts the generalization ability and robustness of the network.
@@
To assess ResNet-50’s ability to generalize to real-world data, we employed Mag1c to invert methane concentration distribution maps derived from RPISMA data in the WENT region and manually identified 28 methane plumes. We estimated their emission rates using both the IME method and the pre-trained ResNet-50 model, constructed scatter plots, and performed linear curve fitting. As shown in Fig. 10, the scatter plot illustrates that the majority points are aligned closely to a straight line, indicating that the predictions of both approaches are typically in good agreement with the real-world dataset. However, it is also apparent that the ResNet-50 model shows a tendency to saturate in areas with lower emission rates.
@@
We manually extracted 28 methane plumes from the inversion results of PRISMA’s WENT region and predicted their emission rates using the IME method and the trained ResNet-50 model separately. The prediction results are shown in the figure. Both methods yielded comparable results. However, it can also be observed that ResNet-50 shows some saturation in regions with low emission rates.
@@ IV-E Serial Network
Although individual models demonstrated high accuracy for each specific task, the efficacy of serializing multi-network models still warrants further examination. Initially, we combined Mask R-CNN and ResNet-50 to handle the tasks of segmenting methane plumes from concentration maps and estimating their emission rates. We assessed the accuracy of these emission rate estimates in the validation set under three scenarios: 1. For true positive segmented plumes, we used pre-trained ResNet-50 to estimate emission rates and compared them with the label’s emission rates. 2. For false-positive segmented plumes, we similarly used ResNet-50 to estimate their emission rates and compared them against 0 kg/h. 3. For misdetected plumes, we assumed their emission rate estimates to be 0 kg/h and compared these estimates with the corresponding labels. Ultimately, RMSE and MAE between all predicted emission rates and label values were calculated (Table IV). The results indicated that the accuracy of the emission rate estimation using Mask R-CNN + ResNet-50 was significantly lower than using ResNet-50 for single-task emission rate estimation. It indicates that the serialization of Mask R-CNN and ResNet-50 could introduce additional errors in the overall performance of methane flux rate estimation.
@@
To investigate additional sources of error, we chose segmented plumes with significant errors (e.g., exceeding 500 kg/h). The analysis identified two additional error sources: 1. Errors stemming from small patches that are essentially noise; 2. Errors arising from false positive patches.
We examined the distribution of minor noise patches and discovered that the majority are defined by a mask size of less than 300 pixels, while authentic plume mask sizes seldom fall below 300 pixels. As a result, we exclude segmented masks smaller than 300 pixels in size and reassess the accuracy of Mask R-CNN + ResNet-50 for estimating emission rates.
@@ IV-F Multi-task Learning I (MTL-01)
We trained the $MTL-01$ network, and after 30 epochs, $MTL-01$ outperformed Mask R-CNN in the plume segmentation task. As a result, the loss in estimating the emission rates of multiple methane sources was also lower than when using a serialized approach with $Mask\,R-CNN$ and $ResNet-50$. Additionally, by simulating a smaller training dataset with both methods, $MTL-01$ demonstrated even higher accuracy compared to the serialized training of $Mask\,R-CNN$ and $MTL-01$. It is evident that the separation of extra plumes is suppressed (Fig. 9), which lowers the predicted emission rate. $MTL-01$ mainly segments the core regions of the jet stream while overlooking some smaller sections of the tail. This has led to improvements in both the plume segmentation task and the emission rate estimation (Table II, Table IV).
Figure 11: Comparison of the sequential network and multi-task learning network I (MTL-01).
@@ IV-G Multi-task Learning II (MTL-02)
We trained the MTL-02 model, which reached convergence after 30 epochs, showing a rapid decline in both loss function values and validation accuracy values. In the end, MTL-02 obtained slightly better accuracy in methane concentration inversion than U-net (Table I), and superior accuracy in plume segmentation compared to Hyper Mask R-CNN (Table II). This supports the notion that multi-task learning could enhance the convergence rate and fitting accuracy of the model, likely due to the shared information among various labels.
@@ IV-H Limitations
Monitoring methane plumes is an intricate issue, and the current research in this domain is neither thoroughly explored nor systematic. Our proposed method also has some constraints: 1. The shapes of the simulated plumes that we used mostly adhere to a standard Gaussian model, which represents an ideal scenario. However, in reality, methane emissions from point sources are influenced by a variety of factors like topography and humidity, leading to non-Gaussian plumes. Whether our method applies to non-Gaussian plumes remains an open question; 2. The maximum overlap ratio of the plumes that we examined is $15\%$. However, naturally, many sources of methane emission are situated close to each other, and the overlap ratio of two plumes can be significantly higher than $15\%$. In such scenarios, even though we can generate highly overlapped methane plumes, it is challenging to provide precise training labels, necessitating the exploration of point-based semi-supervised learning techniques to achieve reasonable segmentation outcomes for these highly overlapped plumes; 3. The acquisition of hyperspectral satellite imagery is significantly affected by weather conditions, and hyperspectral image data sources are scarce in low-latitude coastal regions. Hence, a solution relying solely on hyperspectral images may not fulfill high-frequency monitoring needs. It is essential to devise an ensemble estimation method based on multiple data sources.
@@ V Conclusion
Currently, there are significant uncertainties in monitoring fugitive methane emissions using satellite imaging spectrometers. We have created a demonstrative method utilizing deep learning models and simulated hyperspectral images for methane concentration inversion, plume segmentation, and emission rate estimation in landfills. Specifically, U-net was employed for methane concentration inversion, yielding higher accuracy and reduced noise compared to mag1c. Mask R-CNN was used for plume segmentation, eliminating subjectivity in the process. For estimating methane plume emission rates, ResNet-50 is utilized to bypass the dependence on wind speed measurements. We also addressed additional errors in emission rate estimation caused by serial networks by developing a multi-task model (MTL-01) to reduce these errors and enhance the precision of plume segmentation. Furthermore, a multi-task model (MTL-02) that performs both methane concentration inversion and plume segmentation was proposed, achieving higher validation accuracy than individual single-task models. Finally, we highlighted the limitations of the current approach, emphasizing the need for future research on highly overlapped plumes, non-Gaussian plumes, and the integration of multi-source data for inversion.